A constant torque of 2 kN-m is exerted on a crankshaft to start the engine. The flywheel has a mass of 1800 kg and radius of gyration 1 m. If there is a resisting torque of 1 kN-m find the speed of the engine after 1 minute.


[Hint. Effective- torque = 2-1 = 1 kN-m).


[Ans. 320.9 r.p.m.]


1
Expert's answer
2021-02-18T18:39:59-0500
"(\tau_2-\tau_1)=I\frac{\omega}{t}\\(2-1)1000=(1800)(1)^2\frac{2\pi\omega}{60(60)}\\\omega=318.3\ rpm"